During a basketball game, the Robins scored 17 fewer points than the Cardinals. Together, both teams scored a total of 145 points. How many points did the Robins score during the game?
Answer by RAY100(1637)

Let r= # points Robbins,,,,and c= # points Cardinals
.
r=c-17
.
r+c = 145,,,,,subst
.
(c-17) +c =145
.
2c-17=145
.
2c= 162
.
c = 81
.
r=c-17 = 81-17 = 64
.
check,,,,64+81= 145,,,ok
.
r=c-17,,,,,64=81-17,,,,ok
.